Online chat rooms are dominated by the young. Teens are the biggest users. If we look only at adult internet users (aged 18 and over) 47% of the 18 to 29 age group chat, as do 21% of the 30 to 49 age group and just 7% of those 50 and over. The distribution of age break down for these internet users is: 30% of the adult internet users are 18 to 29, another 50% are 30 to 49 and the remaining 20% are 50 and over. What percentage of adult chat room participants are aged 50 and over?
Answer #1

P(adult chat room participants) = 0.3 * 0.47 + 0.5 * 0.21 + 0.2 * 0.07 = 0.26

P(aged 50 and over | adult room chat participants) = 0.2 * 0.07 / 0.26 = 0.0538 = 5.38%